OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

\[ y - b = m(x - a)\\ y - 6 = -3(x + 1)\\ y - 6 = -3x - 3\\ 3x + y - 3 = 0\]